Determine the measure of angle "a"
I have one corner that is labeled 108* (top right) and the other corner (bottom left) is labeled 162*. The right hand (bottom right) corner is labeled "a"
It is my understanding that this must add up to 180*. But, you have to remember that I really do not understand this at all.

Answer by whisper(27)

The 3 angles of a triangle do add up to 180*.
So we have one that is:
108* and one that is :
163*
Thus the equation is:
108 + 163 + a = 180
275+a = 180
275 - 275 + a = 180 - 275
a = -95
Since this is not possible this cannot be a triangle.
Do you have the correct angle measures?
Please don't hesitate to ask further questions!